# A non-abelian version of Deligne’s Fixed Part Theorem

### v4

@@ 5.3. The Mumford-Tate group of a polarizable variation of ${\mathbb{Q}}$-Hodge structure
To a polarizable ${\mathbb{Q}}$-Hodge structure $(V,\mathsf{F})$ of weight $w$ one associates the Mumford-Tate group ${\rm MT}(V,\mathsf{F})$. It can be characterized as the smallest linear algebraic subgroup of $\mathrm{GL}(V)$ over ${\mathbb{Q}}$ the complex points of which comprise the maps $\phi_{\lambda}\colon V_{\mathbb{C}}\to V_{\mathbb{C}}$ defined by $\lambda^{i}\overline{\lambda}^{j}$ on $\mathsf{F}^{i}\cap\overline{\mathsf{F}}^{j}$, where $i+j=w$. In particular ${\rm MT}(V,\mathsf{F})$ is connected.
@@
The Mumford-Tate group ${\rm MT}(\mathbb{L}_{x},F_{x})$ normalizes the Zariski closure of the monodromy group ${\rm Mo}(\mathbb{L},x)$ in ${\mathrm{GL}}(\mathbb{L}_{x})$.
@@
The Mumford-Tate group ${\rm MT}(V,\mathsf{F})$ and the group ${\rm Mo}(\mathbb{L},x)$ are Tannaka groups in view of the following observation. Let $V$ be a $\mathbb{Q}$-vector space and $G\subset\mathrm{GL}(V_{\mathbb{C}})$ a abstract subgroup. There is an associated Tannaka category $\mathbf{T}$. Its objects are all subquotients of $V^{\otimes n}\otimes(V^{\vee})^{\otimes m}$ preserved over ${\mathbb{C}}$ by the $G$-action and its morphisms are ${\mathbb{Q}}$-linear maps compatible with the $G$-action over ${\mathbb{C}}$. Then the Tannaka group $\mathrm{Tann}(\mathbf{T})$ with respect to the obvious fibre funtor over ${\mathbb{Q}}$ is the same as the smallest linear algebraic subgroup of $\mathrm{GL}(V)$ over ${\mathbb{Q}}$ such that its complexification contains $G$.
@@
Recall that polarizable variations of mixed ${\mathbb{Q}}$-Hodge structure form a Tannaka category for which every point $x\in X$ induces a fibre functor. We think of any polarizable pure variation of ${\mathbb{Q}}$-Hodge structure as an object of this Tannaka category. We denote by ${\rm Tann}((\mathbb{L},{\sf F}),x)\subset{\mathrm{GL}}(\mathbb{L}_{x})$ its Tannaka group with respect to the fibre functor at $x$. Note that ${\rm Mo}(\mathbb{L},x)$ is the Tannaka group of $\mathbb{L}$ in the Tannaka category of ${\mathbb{Q}}$-local systems, and that by Remark 5.9 the Tannaka group of a polarizable $--{\mathbb{Q}}\text{-}$Hodge structure is its Mumford-Tate group. The inclusion of Tannaka groups
@@
As an isomorphism of two fibre functors induces an isomorphism of Tannaka groups this shows that the isomorphism ${\mathrm{GL}}(\mathbb{L}_{x})\cong{\mathrm{GL}}(\mathbb{L}_{y})$, induced by a path between $x,y\in X$, gives us an isomorphism ${\rm MT}((\mathbb{L},{\sf F}),x)\cong{\rm MT}((\mathbb{L},{\sf F}),y)$. In this sense the Mumford-Tate groups at different points of $X$ of a variation of Hodge structure on $X$ form a local system, i.e.  up to isomorphism it is independent of the choice of the base point.
@@ 5.4. Reminder on Simpson’s non-abelian Hodge locus
In this section we formulate a generalization of Simpson’s results in [Sim97, Sec. 12] to good holomorphic maps $f\colon X\to S$ in the sense of Section 1. In contrast to Simpson we formulate the result only “on the Betti side” of the non-abelian Hodge correspondence.
Let $f\colon X\to S$ be a good holomorphic map. As recalled in Subsection 2.2, $f$ is a topological fibre bundle, so the sheaf of pointed sets $R^{1}f_{*}{\mathrm{GL}}_{r}({\mathbb{C}})$ is a local system on $S$, with fibre $H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$, the set of isomorphism classes of rank $r$ complex local systems on $X_{s}$. Its associated étale space is a covering $\epsilon:T\to S$ ([Bou16, I.86 Prop. 9]). We endow $T$ with the usual complex manifold structure. As a topolocial space $T$ can be constructed as
@@
We write $[\mathbb{L}_{s}]\in H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$ for the isomorphism class of a local system $\mathbb{L}_{s}$ on $X_{s}$. Simpson defines the non-abelian Hodge locus or Noether-Lefschetz locus $\mathrm{NL}={\rm NL}(f,r)$ of rank $r$ as the subset of $T$ consisting of those $[\mathbb{L}_{s}]$ such that $\mathbb{L}_{s}$ underlies a ()-$({\mathbb{Z}}){\mathbb{C}}\text{-}{\sf PVHS}$ on $X_{s}$.
@@ 6.1. Setting
Let $f\colon X\to S$ be a good holomorphic map which satisfies one of the conditions I) - III) from Section 1.
@@
In order to deduce Corollary 1.5 one only has to observe that semi-simple representations of surface groups which have finite orbit under the mapping class group, or what is the same finite orbit under the fundamental of the moduli of those Riemann surfaces, have semi-simple algebraic monodromy groups. This follows from Lemma 4.2 and the fact that in rank one such a representation has finite monodromy, see e.g. [BKMS18, Thm. 1.1].
@@
We prove 1) $\Rightarrow$ 3). By Theorem 5.1(v) the algebraic monodromy group of $\mathbb{L}_{s}$ is semi-simple, so we can apply Theorem 1.4 to deduce the existence of an algebraically isomonodromic extension $\mathbb{L}$ of $\mathbb{L}_{s}$, which can be defined over $\mathbb{Z}$. The polarization $Q_{s}$ essentially extends uniquely to a perfect pairing $Q\colon\mathbb{L}\otimes\overline{\mathbb{L}}\to K$. In order to show that the Hodge filtration $\mathsf{F}_{s}$ extends uniquely to a filtration of $\mathbb{L}\otimes_{\mathbb{C}}\mathcal{O}_{X}$ by holomorphic subbundles $\mathsf{F}^{i}$ which make $(\mathbb{L},\mathsf{F},Q)$ a $K\text{-}{\sf PVHS}$ we may assume without loss of generality that $K=\mathbb{C}$. Let
@@
be the canonical decomposition, where $\mathbb{L}_{i}$ are non-isomorphic simple local systems. By Lemma 4.9 we obtain that $\mathbb{L}_{i}|_{X_{s}}$ is simple for all indices $i$, after replacing $S$ by a finite étale covering. Then by Theorem 5.1(iv) there are -${\mathbb{C}}\text{-}{\sf PVHS}$ $(\mathbb{L}_{i}|_{X_{s}},F_{(i),s},Q_{(i),s})$ and -${\mathbb{C}}\text{-}{\sf PHS}$ $(V_{i},F_{(i),V},Q_{(i),V})$ such that $(\mathbb{L}_{s},F_{s},Q_{s})$ is the direct sum of their tensor products via the isomorphism (6.1).
@@
We finally prove 5) $\Rightarrow$ 1), the implication 4) $\Rightarrow$ 1) is analogous and we omit the details. Using Artin approximation [Art69] we can approximate the formal filtration $\hat{\mathsf{F}}$ by a formal filtration with the same fibre over $s$ but which extends to a (relative) Griffiths transversal filtration by subbundles $\mathsf{F}$ of the isomonodromic extension
@@
over $X_{\Delta}$, where $\Delta\subset S$ is a suitable contractible open neighborhood of $s$. Let $\mathbb{L}_{\Delta}$ be the extension of the local system $\mathbb{L}$ to $X_{\Delta}$, let similarly $Q_{\Delta}$ denote the extension of the polarization. As in Appendix B, see Claim A.3, the set of $t\in\Delta$ such that $(\mathbb{L}_{\Delta}|_{X_{t}},\mathsf{F}|_{X_{t}},Q_{\Delta}|_{X_{t}})$ is a -${\mathbb{C}}\text{-}{\sf PVHS}$ is open in $\Delta$, note that in the non-proper case this observation relies on [KU09, Prop. 7.1.1]. By the same reasoning as in the proof of Theorem 3.1 the connected component of the non-abelian Hodge locus $\mathrm{NL}(f,r)$ containing $[\mathbb{L}_{s}]$, see Subsection 5.4, is a finite étale covering of $S$. So the monodromy orbit of $[\mathbb{L}_{s}]$ is finite.
@@ Appendix A Non-abelian Hodge loci (after Simpson)
In this appendix we explain how to generalize Simpson’s proof of Theorem 5.11 from the proper case to good holomophic maps $f$. We assume in the following that $f\colon X\to S$ is good holomorphic, where $S$ can be an arbitrary complex manifold. As the statement is local on $S$, we can fix $s_{0}\in S$ and allow ourselves to replace $S$ by an open neighborhood of $s_{0}$. We assume without loss of generality that $S$ is contractible.
@@
Let us consider the subset $Z\subset G\times_{S}\overline{X}$ consisting of those points $(\xi,x)$ such that either $x\in X$ and $\mathsf{F}_{\xi,x}$ is a Hodge filtration polarized by $Q$ or $x\in\overline{X}\setminus X$ and $\mathsf{F}_{\xi,x}$ together with the nilpotent monodromy operator and $Q$ is a nilpotent orbit. We have to show that $Z$ is an open subset. Indeed this implies that $G^{0}$ is open as it is then equal to the complement of the image of the closed subset $(G\times_{S}\overline{X})\setminus Z$ along the proper projection $G\times_{S}\overline{X}\to G$.
@@
In the proof of Claim A.4 we need the following continuity result, Proposition A.5, for a fibrewise harmonic metric due to T. Mochizuki [Moc09, Prop. 4.2]. We note that $\det\mathbb{M}$ is finite by Theorem 5.1(v), so we can fix a positive flat Hermitian metric $h^{\det}$ on $\det\mathbb{M}$, which is automatically pluri-harmonic, to be concrete take $h^{\det}=\pm Q^{\det}$.
@@
Let $h\colon\mathcal{E}|_{X}\times\overline{\mathcal{E}}|_{X}\to\mathbb{C}$ be the unique fibrewise harmonic, tame and purely imaginary metric with determinant $h^{\det}$. Then $h$ is continuous.
@@
Consider a point $x\in X_{\psi(\xi)}$ around which 1) holds; note that a priori there might be finitely many points for which it fails. The orthogonal complement $(\mathsf{F}^{i}_{\xi,x})^{\perp}$ of $\mathsf{F}^{i}_{\xi,x}$ in $\mathbb{M}_{x}$ with respect to $Q_{x}$ and with respect to the fibrewise harmonic metric $h_{x}$ from Proposition A.5 coincide by continuity since they coincide on $\mathcal{E}_{\psi(\xi(n))}^{i}$ for all $n$ by the description of the metric $h$ as a Hodge metric in terms of $Q$ on the fibre of $X$ over $\psi(\xi(n))$. As $h$ is positive definite we deduce that $\mathbb{M}_{x}=\mathsf{F}^{i}_{\xi,x}\oplus(\mathsf{F}^{i}_{\xi,x})^{\perp}$. In other words $Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}$ is a perfect pairing. For the same reason $(-1)^{i}Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}\cap(\mathsf{F}^{i+1}_{\xi,x})^{\perp}$ is positive definite as by continuity it agrees with $h_{x}$ there. So around $x$ we get a variation of Hodge structure $\mathsf{F}_{\xi}$.
In order to conclude we have to show that the torsion vanishes in the coherent sheaf $\mathcal{E}_{\psi(\xi)}/\mathsf{F}^{i}_{\xi,\psi(\xi)}$ for any $i$. Let $\overline{\xi}(n)$ correspond to the orthogonal Hodge filtration $\overline{\mathsf{F}}^{a}=({\mathsf{F}}^{-a})^{\perp}$ on $X_{\psi(\xi(n))}$ continuously extended to the anti-holomorphic Deligne extension ¯$\overline{\mathcal{E}}_{\psi(\xi(n))}$ on $\overline{X}_{\psi(\xi(n))}$. By the properness of $G$ over $S$, Proposition A.2 1), we can replace $(\xi(n))_{n}$ by a subsequence such that $\overline{\xi}(n)$ converges to some $\overline{\xi}$ in the anti-holomorphic version of $G$. By the continuity of the degree on $G$ and by Lemma 5.5 we obtain
@@
for all $i$ as this holds for the Hodge filtrations corresponding to $\xi(n)$ for all $n$. By Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], the saturations $\tilde{\mathsf{F}}_{\xi,\psi(\xi)}$ of $\mathsf{F}_{\xi,\psi(\xi)}$ and $\tilde{{\mathsf{F}}}_{\overline{\xi},\psi(\xi)}$ of ${\mathsf{F}}_{\overline{\xi},\psi(\xi)}$ form a variation of Hodge structure on $X_{\psi(\xi)}$ of weight $-1$, so
@@
$\deg(\mathsf{F}^{i}_{\xi,\psi(\xi)})\leq\deg(\tilde{\mathsf{F}}^{i}_{\xi,\psi(\xi)}),\quad\deg({\mathsf{F}}^{-i}_{\overline{\xi},\psi(\xi)})\leq\deg(\tilde{{\mathsf{F}}}^{-i}_{\overline{\xi},\psi(\xi)})$

### v5

@@ 5.3. The Mumford-Tate group of a polarizable variation of $\mathbb{Q}$}-Hodge structure
To a polarizable $\mathbb{Q}$}-Hodge structure $(V,\mathsf{F})$ of weight $w$ one associates the Mumford-Tate group ${\rm MT}(V,\mathsf{F})$. It can be characterized as the smallest linear algebraic subgroup of $\mathrm{GL}(V)$ over $\mathbb{Q}$} the complex points of which comprise the maps $\phi_{\lambda}\colon V_{\mathbb{C}}\to V_{\mathbb{C}}$ for $\lambda\in\mathbb{C}$ defined by $\lambda^{i}\overline{\lambda}^{j}$ on $\mathsf{F}^{i}\cap\overline{\mathsf{F}}^{j}$, where $i+j=w$. In particular, ${\rm MT}(V,\mathsf{F})$ is connected.
@@
The Mumford-Tate group ${\rm MT}(\mathbb{L}_{x},F_{x})$ normalizes the Zariski closure ${\rm Mo}(\mathbb{L},x)$ of the monodromy group in ${\mathrm{GL}}(\mathbb{L}_{x})$.
@@
The Mumford-Tate group ${\rm MT}(V,\mathsf{F})$ and the group ${\rm Mo}(\mathbb{L},x)$ are Tannaka groups in view of the following observation. Let $V$ be a $\mathbb{Q}$-vector space and $G\subset\mathrm{GL}(V_{\mathbb{C}})$ a abstract subgroup. There is an associated Tannaka category $\mathbf{T}$. Its objects are all subquotients of $V^{\otimes n}\otimes(V^{\vee})^{\otimes m}$ preserved over $\mathbb{C}$} by the $G$-action and its morphisms are $\mathbb{Q}$}-linear maps compatible with the $G$-action over $\mathbb{C}$}. Then the Tannaka group $\mathrm{Tann}(\mathbf{T})$ with respect to the obvious fiber functor over $\mathbb{Q}$} is the same as the smallest linear algebraic subgroup of $\mathrm{GL}(V)$ over $\mathbb{Q}$} such that its complexification contains $G$.
@@
[…]
@@
As an isomorphism of two fiber functors induces an isomorphism of Tannaka groups this shows that the isomorphism ${\mathrm{GL}}(\mathbb{L}_{x})\cong{\mathrm{GL}}(\mathbb{L}_{y})$, induced by a path between $x,y\in X$, gives us an isomorphism ${\rm MT}((\mathbb{L},{\sf F}),x)\cong{\rm MT}((\mathbb{L},{\sf F}),y)$. In this sense the Mumford-Tate groups at different points of $X$ of a variation of Hodge structure on $X$ form a local system, i.e. up to isomorphism it is independent of the choice of the base point.
@@ 5.4. Reminder on Simpson’s non-abelian Hodge locus
In this section we formulate a generalization of Simpson’s results in [Sim97, Sec. 12]. In contrast to Simpson we formulate the result only “on the Betti side” of the non-abelian Hodge correspondence.
Let $f\colon X\to S$ be a good holomorphic map. As recalled in Subsection 2.2, $f$ is a topological fiber bundle, so the sheaf of pointed sets $R^{1}f_{*}{\mathrm{GL}}_{r}({\mathbb{C}})$ is a local system on $S$, with fiber $H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$, the set of isomorphism classes of rank $r$ complex local systems on $X_{s}$. Its associated étale space is a covering $\epsilon:T\to S$ ([Bou16, I.86 Prop. 9]). We endow $T$ with the usual complex manifold structure. As a topological space $T$ can be constructed as
@@
We write $\mathbb{L}_{s}]\in H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$ for the isomorphism class of a local system $\mathbb{L}_{s}$ on $X_{s}$. Simpson defines the non-abelian Hodge locus or Noether-Lefschetz locus $\mathrm{NL}={\rm NL}(f,r)$ of rank $r$ as the subset of $T$ consisting of those $\mathbb{L}_{s}]$ such that $\mathbb{L}_{s}$ underlies a $({\mathbb{Z}}){\mathbb{C}}\text{-}{\sf PVHS}$ on $X_{s}$.
@@ 6.1. Setting
Let $f\colon X\to S$ be a good algebraic morphism satisfying Ass from Section 1.
@@
In order to deduce Corollary 1.5 one only has to observe that semisimple representations of surface groups which have finite orbit under the mapping class group, or, what is the same, finite orbit under the fundamental group of the moduli of those Riemann surfaces, have semisimple algebraic monodromy groups. This follows from Lemma 4.2 and the fact that in rank one such a representation has finite monodromy, see e.g. [BKMS18, Thm. 1.1].
@@
We prove 1) $\Rightarrow$ 3). By Theorem 5.1(v) the algebraic monodromy group of $\mathbb{L}_{s}$ is semisimple, so we can apply Theorem 1.4 to deduce the existence of an algebraically isomonodromic extension $\mathbb{L}$ of $\mathbb{L}_{s}$, which can be defined over $\mathbb{Z}$. The polarization $Q_{s}$ essentially extends uniquely to a perfect pairing $Q\colon\mathbb{L}\otimes\overline{\mathbb{L}}\to K$. In order to show that the Hodge filtration $\mathsf{F}_{s}$ extends uniquely to a filtration of $\mathbb{L}\otimes_{\mathbb{C}}\mathcal{O}_{X}$ by holomorphic subbundles $\mathsf{F}^{i}$ which make $(\mathbb{L},\mathsf{F},Q)$ a $K\text{-}{\sf PVHS}$ we may assume without loss of generality that $K=\mathbb{C}$. Let
@@
be the canonical decomposition, where $\mathbb{L}_{i}$ are non-isomorphic simple local systems and where $V_{i}=\mathrm{Hom}(\mathbb{L}_{i},\mathbb{L})$ are complex vector spaces. By Lemma 4.9 we can assume that $\mathbb{L}_{i}|_{X_{s}}$ is simple for all indices $i$, after replacing $S$ by a finite étale covering. Then by Theorem 5.1(iv) there are -${\mathbb{C}}\text{-}{\sf PVHS}$ $(\mathbb{L}_{i}|_{X_{s}},F_{(i),s},Q_{(i),s})$ and -${\mathbb{C}}\text{-}{\sf PHS}$ $(V_{i},F_{(i),V},Q_{(i),V})$ such that $(\mathbb{L}_{s},F_{s},Q_{s})$ is the direct sum of their tensor products via the isomorphism (6.1).
@@
We finally prove 5) $\Rightarrow$ 1), the implication 4) $\Rightarrow$ 1) is analogous and we omit the details. Using Artin approximation [Art69] we can approximate the formal filtration $\hat{\mathsf{F}}$ by a formal filtration with the same fiber over $s$ but which extends to a (relative) Griffiths transversal filtration by subbundles $\mathsf{F}$ of the isomonodromic extension
@@
over $X_{\Delta}$, where $\Delta\subset S$ is a suitable contractible open neighborhood of $s$. Let $\mathbb{L}_{\Delta}$ be the extension of the local system $\mathbb{L}$ to $X_{\Delta}$, let similarly $Q_{\Delta}$ denote the extension of the polarization. As in Appendix B, see Claim A.3, the set of $t\in\Delta$ such that $(\mathbb{L}_{\Delta}|_{X_{t}},\mathsf{F}|_{X_{t}},Q_{\Delta}|_{X_{t}})$ is a -${\mathbb{C}}\text{-}{\sf PVHS}$ is open in $\Delta$, note that in the non-proper case this observation relies on [KU09, Prop. 7.1.1]. By the same reasoning as in the proof of Theorem 3.1 (5) $\Rightarrow(1)$, the connected component of the non-abelian Hodge locus $\mathrm{NL}(f,r)$ containing $\mathbb{L}_{s}]$, see Subsection 5.4, is a finite étale covering of $S$. So the monodromy orbit of $\mathbb{L}_{s}]$ is finite.
@@ Appendix A Non-abelian Hodge loci (after Simpson)
In this appendix we explain how to generalize Simpson’s proof of Theorem 5.11 from the proper case to good holomorphic maps $f$. We assume in the following that $f\colon X\to S$ is good holomorphic, where $S$ can be an arbitrary complex manifold. As the statement is local on $S$, we can fix $s_{0}\in S$ and allow ourselves to replace $S$ by an open neighborhood of $s_{0}$. We assume without loss of generality that $S$ is contractible.
@@
Let us consider the subset $Z\subset G\times_{S}\overline{X}$ consisting of those points $(\xi,x)$ such that either $x\in X$ and $\mathsf{F}_{\xi,x}$ is a Hodge filtration polarized by $Q$ or $x\in\overline{X}\setminus X$ and $\mathsf{F}_{\xi,x}$ together with the nilpotent monodromy operator and $Q$ is a nilpotent orbit. We have to show that $Z$ is an open subset. Indeed, this implies that $G^{0}$ is open as it is then equal to the complement of the image of the closed subset $(G\times_{S}\overline{X})\setminus Z$ along the proper projection $G\times_{S}\overline{X}\to G$.
@@
In the proof of Claim A.4 we need the following continuity result, Proposition A.5, for a fiberwise harmonic metric due to T. Mochizuki [Moc09, Prop. 4.2]. We note that $\det\mathbb{M}$ is finite by Theorem 5.1(v), so we can fix a positive flat hermitian metric $h^{\det}$ on $\det\mathbb{M}$, which is automatically pluri-harmonic, to be concrete take $h^{\det}=\pm Q^{\det}$.
@@
Let $h\colon\mathcal{E}|_{X}\times\overline{\mathcal{E}}|_{X}\to\mathbb{C}$ be the unique fiberwise harmonic, tame and purely imaginary metric with determinant $h^{\det}$. Then $h$ is continuous.
@@
Consider a point $x\in X_{\psi(\xi)}$ around which 1) holds; note that a priori there might be finitely many points for which it fails. The orthogonal complement $(\mathsf{F}^{i}_{\xi,x})^{\perp}$ of $\mathsf{F}^{i}_{\xi,x}$ in $\mathbb{M}_{x}$ with respect to $Q_{x}$ and with respect to the fiberwise harmonic metric $h_{x}$ from Proposition A.5 coincide by continuity since they coincide on $\mathcal{E}_{\psi(\xi(n))}^{i}$ for all $n$ by the description of the metric $h$ as a Hodge metric in terms of $Q$ on the fiber of $X$ over $\psi(\xi(n))$. As $h$ is positive definite we deduce that $\mathbb{M}_{x}=\mathsf{F}^{i}_{\xi,x}\oplus(\mathsf{F}^{i}_{\xi,x})^{\perp}$. In other words $Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}$ is a perfect pairing. For the same reason $(-1)^{i}Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}\cap(\mathsf{F}^{i+1}_{\xi,x})^{\perp}$ is positive definite as by continuity it agrees with $h_{x}$ there. So around $x$ we get a variation of Hodge structure $\mathsf{F}_{\xi}$.
In order to conclude we have to show that the torsion vanishes in the coherent sheaf $\mathcal{E}_{\psi(\xi)}/\mathsf{F}^{i}_{\xi,\psi(\xi)}$ for any $i$. Let $\overline{\xi}(n)$ correspond to the orthogonal Hodge filtration $\overline{\mathsf{F}}^{a}=({\mathsf{F}}^{-a})^{\perp}$ on $X_{\psi(\xi(n))}$ continuously extended to the anti-holomorphic Deligne extension $\overline{\mathcal{E}}_{\psi(\xi(n))}$ on $\overline{X}_{\psi(\xi(n))}$. By the properness of $G$ over $S$, Proposition A.2 1), we can replace $(\xi(n))_{n}$ by a subsequence such that $\overline{\xi}(n)$ converges to some $\overline{\xi}$ in the anti-holomorphic version of $G$. By the continuity of the degree on $G$ and by Lemma 5.5 we obtain
@@
for all $i$ as this holds for the Hodge filtrations corresponding to $\xi(n)$ for all $n$. By Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], the saturations $\tilde{\mathsf{F}}_{\xi,\psi(\xi)}$ of $\mathsf{F}_{\xi,\psi(\xi)}$ and $\tilde{{\mathsf{F}}}_{\overline{\xi},\psi(\xi)}$ of ${\mathsf{F}}_{\overline{\xi},\psi(\xi)}$ form a variation of Hodge structure on $X_{\psi(\xi)}$ of weight 1-1, so
@@
[…]